An open cylindrical tank of 20 m high and 30 m in diameter is filled with water. Given





that the atmospheric pressure,





P





=1.013 × 105 N/m2





.





i) Determine the water pressure on the bottom of the tank.





ii) Calculate the total force on the bottom of tank.






1
Expert's answer
2022-02-09T06:26:24-0500

Given:

"h=\rm 20\:\rm m"

"d=\rm 30\:\rm m"

"p_a=\rm1.013 *10^5 \:N/m^2"

"\rho=\rm 1000\: kg/m^3"


(i) the water pressure on the bottom of the tank

"p=p_a+g\rho h\\
=\rm1.013 *10^5 +9.8*1000*20=2.97*10^5\:\rm Pa"

(ii) the total force on the bottom of tank

"F=pA=2.97*10^5*3.14*15^2=2.1*10^8\:\rm N"